OpenStudy (anonymous):

Multiply and find the product (x^5+y^2)(x^5-y^2)

OpenStudy (ajprincess):

\((a-b)(a+b)=a^2-b^2\) Plug in \(x^5\) in place of a and \(y^2\) in place of b. Can u do @cdelomas?

OpenStudy (anonymous):

can you help me

OpenStudy (anonymous):

would it be 3

OpenStudy (ajprincess):

no In \(a^2-b^2\) plug in \(x^5\) for a and \(y^2\) for b. nw can u do.

OpenStudy (anonymous):

is it 5*2

OpenStudy (anonymous):

are you still there

OpenStudy (ajprincess):

Nope. \((x^5+y^2)(x^5-y^2)=(x^5)^2-(y^2)^2\) using \(\large(c^d)^e=c^{d*e}\) \((x^5)^2-(y^2)^2=(x^{5*2})-(y^{2*2})\) =\(x^{10}-y^4\) getting this? If u dnt get any of this plz ask:)

OpenStudy (anonymous):

where did c,d,e comefr om

OpenStudy (ajprincess):

They are jst part of general formula>

OpenStudy (anonymous):

iGOTEVe ry thing else

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

So what would beth e product

OpenStudy (anonymous):

hello

OpenStudy (ajprincess):

That's what I gave u. Guess u still didnt get me. so please if u dnt understand anythng ask me. I will try my best to explain u.:)

OpenStudy (anonymous):

no I got it just did not k now if that w as the product.

OpenStudy (anonymous):

thanks

OpenStudy (ajprincess):

welcome:)